The expression for the length of a rectangular garden is x+3 and the expression for width is x-4 if the area of the garden is 60 square feet find the value of x
Answer by stanbon(75887)

The expression for the length of a rectangular garden is x+3 and the expression for width is x-4 if the area of the garden is 60 square feet find the value of x
--------------------------
Area = length*width
----
(x+3)(x-4) = 60
------
x^2 -x - 12 = 60
------
x^2 -x = 72 = 0
------
Factor:
(x-9)(x+x) = 0
---
Positive solution:
x = 9
----
length = x+3 = 12
width = x-4 = 5
---------------------
Cheers,
Stan H.
====================